OpenStudy (anonymous):

Find each percent of change and round to the nearest tenth. 36 to 30

OpenStudy (anonymous):

40

OpenStudy (anonymous):

or else..............

OpenStudy (anonymous):

@all members i got 70 medals

Directrix (directrix):

(36 - 30)/36 = 6/36 = .1667 = 16.67 decrease

OpenStudy (anonymous):

% change = (new value - old value)/old value = (30-36)/36 = -16.6667%=-17%

OpenStudy (anonymous):

It should be -16.7 as it is corrected to the nearest tenth

OpenStudy (jagatuba):

@sarah224 You have now posted 12 questions all of the exact same nature. I have seen some very good explanations with very little feedback from you, (e.i. I understand, I still don't get it) yet you continue to post the same questions. Am I to assume that ese are your homework questions and you are just trying to get others to answer them for you? This is against OpenStudy's anti-cheating policy ( http://openstudy.com/terms-and-conditions ). I know you are new here which is why I'm giving you a heads up instead of an official warning.